OpenStudy (anonymous):

How much force is needed to stop a 100-kg football player if he decelerates at 20 m/s²? A. 20 N B. 2000 N C. 200 N D. 5 N

OpenStudy (anonymous):

Formula for force is F = ma You have 100kg as mass, and acceleration is 20m/s^2 Therefore F = 100*20 = 2000 answer is B

OpenStudy (anonymous):

Thank you so much!